19, 9, 10, 13, 23, ?, 10, 15, 9
Answer by Edwin McCravy(20056)

19, 9, 10, 13, 23, ?, 10, 15, 9
Rule:
1. For 19, 20, 21, ..., write the integer as a product of prime numbers.
2. Add all the prime numbers in the prime factorization.
3. If a number is prime just write that prime as its prime factorization.

The prime factorization of 19 is 19
The prime factorization of 20 is 2*2*5 and 2+2+5 = 9
The prime factorization of 21 is 3*7 and 3+7 = 10
The prime factorization of 22 is 2*11 and 2+11 = 13
The prime factorization of 23 is 23
The prime factorization of 24 is 2*2*2*3 and 2+2+2+3 = 9
The prime factorization of 25 is 5*5 and 5+5 = 10
The prime factorization of 26 is 2*13 and 2+13 = 15
The prime factorization of 27 is 3*3*3 and 3+3+3 = 9

Edwin